# An instrument to link global positioning to the Universe - Observing GNSS satellites with the Australian VLBI array

### v2

@@ 1 Introduction
Accurate timing, navigation and positioning information underpins modern society (GGCE, 2024). It is provided through Global Navigation Satellite Systems (GNSS), such as the Global Positioning System (GPS). Satellite navigation itself relies on accurate orbit information. For this, one needs to disentangle the movement of the satellite in the Earth’s gravity field and the variable rotation of the Earth. Precise Earth orientation parameters are provided through the International Earth Rotation and Reference System Service (IERS), which combines measurements from the Global Geodetic Observing System (GGOS), such as those provided by GNSS or Very Long Baseline Interferometry (VLBI). In VLBI (i.e. Nothnagel et al., 2017; Sovers et al., 1998), a global network of radio telescopes receives signals emitted from extragalactic radio sources, allowing for precise measurement of the distance between the receiving stations (baselines) as well as the orientation of the Earth with respect to a celestial reference frame (ICRF; Charlot et al., 2020). DUT1, the difference between the actual rotation of the Earth and universal time maintained by a set of atomic clocks, is the parameter that is uniquely measured by the VLBI technique.
@@
The consequences of this newly developed technique are twofold: A direct link between VLBI and GNSS will be possible, initially over Australia, and globally if this technique can be transferred to other VGOS telescopes worldwide; presenting a new opportunity with significant implications for an improved TRF. Secondly, this proof-of-concept is just the important piece missing in order to make the VLBI part of the Genesis mission a success: after the mission being granted it has become obvious that the VLBI observations to Genesis are completely novel and in order to be successful, a multitude of technological challenges must be overcome. This becomes achievable with actual field test.
@@ 1.1 State of the art
VLBI observations to spacecrafts are routinely included in (outer) space missions, leveraging on the high precision measurements for orbit determination (Duev et al., 2012; Hanada et al., 2008; Lebreton et al., 2005). When it comes to geodetic VLBI though, incorporation of novel observations is not straightforward. Observing the Chinese Chang’e3 lunar lander with the IVS network (Klopotek et al., 2019) and VLBI observations from the AuScope array to the APOD mission (Sun et al., 2018; Hellerschmied et al., 2018), revealed some of these issues. One of the conclusions from these studies was that optimisation of the technique will require more observations. Böhm & Wolf (2024) nicely summarise current work towards VLBI transmitters on satellites. Multiple groups have studied this technique using simulations (Plank et al., 2014; Anderson et al., 2018; Klopotek et al., 2020; Wolf & Böhm, 2023; Pollet et al., 2023; Schunck et al., 2024a, b) with studies focusing on the technical realisation by Jaradat et al. (2021), Plank et al. (2017), and Schunck et al. (2024). A slightly different observing concept using new hardware is discussed by Skeens et al. (2023).
@@ 2.2 Observations
For the test observations, three satellites of the GPS system (block IIF) were observed. The AuScope array has QRFH VGOS feed horns (Akgiray et al., 2013), with a nominal operating frequency range between 2-14 GHz. The key discovery for this work was that despite the nominal cut-off, GPS signals in L-band were detected with VLBI. This is due to the fact that GPS signals are tremendously strong. For these tests, we typically used blocks of a few hours on either 2 or 3 telescopes. The preparation of these experiments required is minimal (1-2 hrs). Observing itself is done remotely, with the current bottleneck being the data transport to the correlator in Hobart; while Yarragadee data can be e-transferred at good speeds, the data from Katherine needs to be physically shipped until the high speed data connection will become available during 2025. For future full scale sessions (24-hrs) we would expect turnaround times of about 2-3 weeks, which is common for IVS observations.
@@ 2.3.1 Multi-frequency:
Despite using out-of-frequency receivers, GPS signals on the three main frequencies, L1, L2 and L5, could be observed. These out-of-band signals are strongly attenuated by the receivers, which proves beneficial for the subsequent VLBI processing. In previous observations by Plank et al. (2017), the strong signals partially caused saturation or at least problems with the automated gain control systems in the digitization stage. We thus expect better VLBI performance for the new observations. Successful observations were also done to satellites of the Galileo (E5a, E5b, E6 and E1) and Beidou (B2a, B2b, B3 and B1) system at sky frequencies of 1176.45, 1207.14, 1268.52, and 1575.42 MHz respectively. In general, we expect compatibility with all GNSS systems.
@@ 2.3.3 Geodetic observing and processing pipeline:
In contrast to test observations, for bulk data production one has to make sure that all observing and processing goes through standard geodetic/IVS procedures and software. With these tests, we have made new steps towards full incorporation into the geodetic pipeline. An observing plan can be created using state-of-the-art scheduling software, i.e. VieSched++ (Schartner & Böhm, 2019) which allows for the option of satellite targets. As the .skd format (which is typically used for observing) does not support satellite targets yet, we create so-called snap-files (.snp). These can be loaded into the antenna field systems (NASA field system [^2] https://github.com/nvi-inc/fs) that is used to coordinate the observations including antenna steering and recording. Since the AuScope telescopes are not set up for continuous satellite tracking via the field system, we use the satellite command (in the field system), implementing step-wise tracking. For correlation using DiFX (Deller et al., 2007, 2011), the a priori delay model is created through difxcalc via externally created SPICE Kernel files (Acton, 1996; Acton et al., 2018). While the correlation process delivers visibilities, the actual VLBI delay is formed in a process typically referred to as fringe-fitting using the fourfit program as part of the HOPS (MIT/Haystack, 2024) pipeline. Here some minor tweaks had to be made to pass validation checks when using 8-bit sampled data.
@@ 3 Results
A first attempt for analysis of the data was done using the VieVS (Böhm et al., 2018) environment, which has a satellite model implemented. At this stage, however, the theoretical models do not agree well with the observations. We find residuals at the level of tens of nanoseconds. While this is not sufficient for meaningful geodetic analysis yet, we want to point out that the residuals in VieVS are one order of magnitude better than the residuals to the correlator model using difxcalc. It shall also be noted that the residuals show a clear signal rather than noise, which is highly suggestive of a reference frame issue for the delays. This data can now be used for further testing and refinements of the delay modelling in various analysis software.

### v3

# An experiment to observe GNSS signals with the Australian VGOS array

@@ 1 Introduction
Accurate timing, navigation and positioning information underpins modern society (GGCE, 2024). It is provided through Global Navigation Satellite Systems (GNSS), such as the Global Positioning System (GPS). Satellite navigation itself relies on accurate orbit information. For this, one needs to disentangle the movement of the satellite in the Earth’s gravity field and the variable rotation of the Earth. Precise Earth orientation parameters are provided through the International Earth Rotation and Reference System Service (IERS), which combines measurements from the Global Geodetic Observing System (GGOS), such as those provided by GNSS or Very Long Baseline Interferometry (VLBI). In VLBI (i.e. Nothnagel et al., 2017; Sovers et al., 1998), a global network of radio telescopes receives signals emitted from extragalactic radio sources, allowing for precise measurement of the distance between the receiving stations (baselines) as well as the orientation of the Earth with respect to a celestial reference frame (ICRF; Charlot et al., 2020). UT1-UTC, the difference between the actual rotation of the Earth and universal time maintained by a set of atomic clocks, is the parameter that is uniquely measured by the VLBI technique.
@@
The consequences of this newly developed technique are twofold: A direct link between VLBI and GNSS will be possible, initially over Australia, and globally if this technique can be transferred to other VGOS telescopes worldwide; presenting a new opportunity with significant implications for an improved TRF. Secondly, this proof-of-concept is just the important piece missing in order to make the VLBI part of the Genesis mission a success: after the mission being granted it has become obvious that the VLBI observations to Genesis are completely novel and in order to be successful, a multitude of technological challenges must be overcome. This becomes achievable with actual field tests.
@@ 1.1 State of the art
VLBI observations to spacecrafts are routinely included in (outer) space missions, leveraging on the high precision measurements for orbit determination (Duev et al., 2012; Hanada et al., 2008; Lebreton et al., 2005). When it comes to geodetic VLBI though, incorporation of novel observations is not straightforward. Observing the Chinese Chang’e3 lunar lander with the IVS network (Klopotek et al., 2019) and VLBI observations from the AuScope array to the APOD mission (Sun et al., 2018; Hellerschmied et al., 2018), revealed some of these issues. One of the conclusions from these studies was that optimisation of the technique will require more observations. Böhm & Wolf (2024) nicely summarise current work towards VLBI transmitters on satellites. Multiple groups have studied this technique using simulations (Plank et al., 2014; Anderson et al., 2018; Klopotek et al., 2020; Wolf & Böhm, 2023; Pollet et al., 2023; Schunck et al., 2024, 2024) with studies focusing on the technical realisation by Jaradat et al. (2021), Plank et al. (2017), and Schunck et al. (2024). A slightly different observing concept using new hardware is discussed by Skeens et al. (2023).
@@ 2.2 Observations
For the test observations, three satellites of the GPS system (block IIF) were observed. The AuScope array has QRFH VGOS feed horns (Akgiray et al., 2013), with a nominal operating frequency range between 2-14 GHz. The key discovery for this work was that despite the nominal cut-off, GPS signals in L-band were detected with VLBI. This is due to the fact that GPS signals are tremendously strong. After the feed horn and standard VGOS cryogenic low noise amplifiers (LNA) the signal is split. The satellite signal passes through a second-stage LNA and filter stage (500-1650 MHz) and travels via coaxial cable without down-conversion to the control room. The observed frequencies lie out of range for the normal calibration techniques in place at the VGOS telescopes. The approximate sensitivity of our receiving system is on the order of several 100.000 Jy in source equivalent flux density (SEFD).
For these tests, we typically used blocks of a few hours on either 2 or 3 telescopes. The preparation of these experiments required is minimal (1-2 hrs). Observing itself is done remotely, with the current bottleneck being the data transport to the correlator in Hobart; while Yarragadee data can be e-transferred at good speeds, the data from Katherine needs to be physically shipped until the high speed data connection will become available during 2025. For future full scale sessions (24-hrs) we would expect turnaround times of about 2-3 weeks, which is common for IVS observations.
VGOS feed horns are linearly polarised. This means, that we record the circularly polarised signal from the satellite in two polarisations at each station. During correlation, the two data streams from each station are cross-correlated, leading to four correlation products, i.e. XX, XY, YX, and YY. Similar to the approach of Plank et al. (2017), all subsequent investigations are using the XX product only. A thorough combination and calibration for the full polarisation product could be subject of future research. Using all data streams will certainly increase the sensitivity of the system, which is not a limiting factor at the moment. A more serious issue would be if delay offsets were found between the various products, which is yet to be investigated in detail.
The main experiment investigated below is ASO304, observed on October 30, 2024. Figure 2 shows the geometry of this experiment. The observations are split into individual scans, integrating over the scan duration of 20 seconds. Three GPS block IIF satellites (G10, G27, G32) were observed, with a minimal repeat time of the same satellite of 7 minutes.
Figure 2: Geometry of experiment ASO304. The satellite ground tracks are shown for the three observed satellites along with topocentric view angles at the observing stations.
@@ 2.3.1 Multi-frequency:
Despite using out-of-frequency receivers, GPS signals on the three main frequencies, L1, L2 and L5, could be observed. These out-of-band signals are strongly attenuated by the receivers, which proves beneficial for the subsequent VLBI processing. Indeed, these observations do not cause significant variations in the total power received at the first-stage amplifiers or at various stages in the receiving system. This is in stark contrast to previous observations by Plank et al. (2017), where strong signals partially caused saturation or at least problems with the automated gain control systems in the digitization stage. We thus expect better VLBI performance for the new observations. Successful observations were also done to satellites of the Galileo (E5a, E5b, E6 and E1) and Beidou (B2a, B2b, B3 and B1) system at sky frequencies of 1176.45, 1207.14, 1268.52, and 1575.42 MHz respectively. In general, we expect compatibility with all GNSS systems.
@@ 2.3.3 Geodetic observing and processing pipeline:
In contrast to test observations, for bulk data production one has to make sure that all observing and processing goes through standard geodetic/IVS procedures and software. With these tests, we have made new steps towards full incorporation into the geodetic pipeline. An observing plan can be created using state-of-the-art scheduling software, i.e. VieSched++ (Schartner & Böhm, 2019) which allows for the option of satellite targets. As the .skd format (which is typically used for observing) does not support satellite targets yet, we create so-called snap-files (.snp). These can be loaded into the antenna field systems (NASA field system [^2] https://github.com/nvi-inc/fs) that is used to coordinate the observations including antenna steering and recording. Since the AuScope telescopes are not set up for continuous satellite tracking via the field system, we use the satellite command (in the field system), implementing step-wise tracking. An update interval of 5 seconds was found to be sufficient for the GNSS satellites to stay within the main beam of the telescopes. For the conducted experiments, a maximal angular change of $0.046^{\circ}$ per 5 second update interval was found. This is one magnitude lower than the half power beam width of $0.91^{\circ}$ calculated for the 12 m telescopes at L1.
For correlation using DiFX (Deller et al., 2007, 2011), the a priori delay model is created through difxcalc via externally created SPICE Kernel files (Acton, 1996; Acton et al., 2018). While the correlation process delivers visibilities, the actual VLBI delay is formed in a process typically referred to as fringe-fitting using the fourfit program as part of the HOPS (MIT/Haystack, 2024) pipeline. Here some minor tweaks had to be made to pass validation checks when using 8-bit sampled data.
@@
It shall further be noted, that some observations did not lead to a reliable fringe detection and delay measurement. Those were mostly related to satellites of the block III type, with a different signal modulation. Most likely, the characteristic double-peak related to the signal modulation is the reason for unsuccessful delay derivation, though it is also possible that the way the fringe fitting is applied or the lack of proper polarisation handling are causing the issues here. It does emphasize that further research is necessary to account for a variety of GNSS signals.
@@ 3 Results
A first attempt for analysis of the data was done using the VieVS (Böhm et al., 2018) environment, which has a satellite model implemented. Comparing the observed and computed delays we find residuals at the level of a few nanoseconds. The residuals are at a similar level as reported in Plank et al. (2017), though this time there are observations on three baselines compared to their single baseline observations. The residuals are one order of magnitude better to the correlator model using difxcalc. Once more observational data is available, we expect the analysis to provide meaningful geodetic results. The current residuals are thought to mainly be due to uncalibrated clock errors, the troposphere as well as refinements in the delay model that is implemented in the software. In addition, as pointed out throughout this paper, there is yet more work to be done to further understand the VLBI processing (i.e. correlation and fringe fitting process) and delay generation of this unusual signal.